2. At the beginning of 2016, Robotics Inc. acquired a manufacturing facility for $12.6 million. $9.6 million of the purchase price was allocated to the building. Depreciation for 2016 and 2017 was calculated using the straight-line method, a 25-year useful life, and a $1.6 million residual value. In 2018, the estimates of useful life and residual value were changed to 20 total years and $560,000, respectively. What is the depreciation on the building for 2018? (Round answer to the nearest whole dollar.)

Depreciation = ____________________
Answer #1

Depreciation on building = (cost-salvage value) /life

2016 & 2017 = (9600000 - 1600000) / 25 = $320000 per year

Net book value at the end of 2017 = 9600000 - 320000*2 = $8960000

Depreciation for 2018 = (8960000-560000) /18 = $466,667